OpenStudy (anonymous):

2.22x10 to the 8th power x 4.00x10 to the 6th power

OpenStudy (unklerhaukus):

\[2.22\times10^8\times 4.00\times10^6\] \[=2.22\times 4.00\times10^8\times10^6\] \[=8.88\times10^{8+6}\] \[=\quad \dots\]